A train traveling at 7.55 m/s accelerates to a distance of 975 m in 31.5 s. What is the final velocity of the train?
1
Expert's answer
2019-04-25T09:19:22-0400

We can find the final velocity of the train from the kinematic equation:


"d = \dfrac{v_i + v_f}{2} \cdot t,"

here, "v_i = 7.55 m\/s" is the initial velocity of the train, "v_f" is the final velocity of the train, "d = 975 m" is the distance traveled by the train, "t = 31.5 s" is the time during which the train accelerates.

Then, from this formula we can find the final velocity of the train:


"v_f = \dfrac{2d}{t} - v_i = \dfrac{2 \cdot 975 m}{31.5 s} - 7.55 \dfrac{m}{s} = 54.35 \dfrac{m}{s}."

Answer:

"v_f = 54.35 \dfrac{m}{s}."